94). Perimeter of a rectangle. The width of a rectangular playground
is 2x - 5 feet, and the length is 3x+9 feet. Write a
polynomial P(x) that represents the perimeter and then
evaluate this perimeter polynomial if x is 4 feet.

Answer by checkley75(3666)

p(x)=2(2x+5)+2(3x+9)
p(4)=2(2*4+5)+2(12+9)
p(4)=2(8+5)+2*21
p(4)=2*13+42
p(4)=26+42
p(4)=68 ans.